OpenStudy (anonymous):

Heyy guys quick question but for the partial fractions method of integration, can A and B be the same number? YES or NO Thanks :)

OpenStudy (anonymous):

yes.. they can be!

OpenStudy (anonymous):

ex \[\frac{2x}{x^2-1} = \frac{1}{x-1} + \frac{1}{x+1}\] See both A and B = 1 here!

OpenStudy (anonymous):

yessurr my man bless u